# Towards Scalable Automated Alignment of LLMs: A Survey

### v2

@@ 3.2.3 Discussion
Task decomposition and self-play both necessitate LLMs functioning as agents. However, the challenge of aligning LLMs as agents is more complex compared to aligning them as chatbots, as it requires considerations of behavior-level alignment (Pan et al., 2023), the dynamics of environment and self-constraints (Garrabrant and Demski, 2018; Shavit et al., 2023; Yang et al., 2024f). We emphasize the significance of this research direction and advocate for increased efforts in this domain. Furthermore, an even more intricate problem lies in proving the theoretical safety and trustworthiness of multi-agent systems. Although research in this area is nascent (Yang and Wang, 2020; DiGiovanni and Zell, 2021), advancements in game theory (Hazra and Anjaria, 2022), automated theorem proving techniques (Polu and Sutskever, 2020) and real-world simulation technology (Brooks et al., 2024) may offer insights into addressing this challenge.
While adversarial self-play offers a potential solution to the scalable oversight challenge, it can also give rise to the emergence of more deceptive (Hubinger et al., 2024), persuasive and autonomous agents (Tao et al., 2024). Such developments could have significant social impacts and ethical risks, such as the potential for models to generate more persuasive articles than humans, which could be exploited for political manipulation. Encouragingly, several prominent model engine providers have taken steps to monitor and mitigate these potential side effects. For instance, OpenAI’s Preparedness team has established benchmarks related to persuasion and autonomy (OpenAI, 2023a). They categorize model risks into four levels: low, medium, high, and critical, stipulating that models above the high-risk threshold cannot be developed, and those above medium risk cannot be deployed. We call upon the academic community and third-party organizations to dedicate additional efforts to the development and review of a robust safety framework for highly capable agents.
@@ 5.1.3 Reinforcement Learning from AI Feedback
With the improved capabilities of LLMs, directly using them to rank multiple responses can provide preference data (Tunstall et al., 2023; Hong et al., 2023; Guo et al., 2024a; Pace et al., 2024; Yuan et al., 2024b). This ranked preference data can also be produced with minimal human supervision, such as through human-defined principles (Bai et al., 2022b; Sun et al., 2023c) or rules (Kim et al., 2023). To improve the quality of generated preference data, Shi et al. (2024) propose a carefully designed pipeline including reversed instruction tuning, instruction induction, and expert model evaluation.  Liu et al. (2024a) propose to score the response using contrastive prompt pairs as input, which can achieve better performance compared to directly generated feedback with a single prompt.
@@ 5.2 Binary Verifier
Even if the final answer is correct, there may still be errors in the reasoning process, limiting the effectiveness of the outcome verifier. To address this problem, the process verifier is employed to assess the correctness of reasoning step for more detailed verification (Lightman et al., 2023; Uesato et al., 2022). The process verifier can be used to train a more effective reasoner (Ying et al., 2024; Shao et al., 2024). Inspired by the human reasoning mechanism, Zhu et al. (2023b) propose Cooperative Reasoning (CoRe) to produce synthesized training data for reasoning where process verifier is used to generate the feedback of the generation of generator. Many studies are dedicated to training the verifier using automatically generated data due to the difficulty of collecting step-wise supervised signals.  Wang et al. (2024e) and Wang et al. (2024h) train the process verifier with automatically constructed data that is collected by Monte Carlo Sampling. Moreover, the process verifier can be applied in decoding to select the correct reasoning path (Khalifa et al., 2023). Some studies focus on how to complete the final reasoning path efficiently. Ma et al. (2023b) propose a heuristic greedy search algorithm based on the verifier’s feedback. Li et al. (2023c) propose using the verifier to filter the reasoning steps generated using diverse prompts.
@@ 6.2 Human Shared-Values
In order to achieve consensus on what basic principles AI systems should follow among groups of people and reach human shared-values by synthesizing various human ideas into a set of target for model alignment, follow-up efforts are dedicated to enable broader public to collectively shape the behavior of AI systems. For example, in a program titled Democratic inputs to AI444openai.com/blog/democratic-inputs-to-ai, an open call for the design of prototype systems to promote the democratic process of AI was issued. Specifically, it addresses enabling a representative cross-section of people to exchange views and ultimately reach a consensus on the rules AI systems should follow. Subsequently, in the Collective Constitutional AI (CCAI) project (Anthropic, 2023; Huang et al., 2024), researchers establish a multi-stage process for sourcing and integrating public input into LMs based on questionnaire. Through the way of collective constitution drafting, participants can assess the existing constitutional principles, rate their acceptance scores, and propose new constitutions they believe AI should be better to adhere to, thus reduce the bias of aligned LLMs. Beyond these pioneering research, large language models themselves are also involved to help conclude shared-values from human crowds more efficiently. Klingefjord et al. (2024) propose Moral Graph Elicitation (MGE) to elicit and reconcile values from human participants, which utilizes an LLM to interview participants about their values in particular contexts.
@@ 7 Underlying Mechanism of Automated Alignment
  • RQ2: Why does Self-feedback work? Self-feedback is widely applied across various paradigms of automated alignment, such as serving as an inductive bias in section 3.1.2, constructing data for preference learning in section 5 and acting as a crucial technique for automated iterative alignment (Yuan et al., 2024b; etc.). Understanding why does it work is vital for comprehending and enhancing all paradigms involving it.
@@ 7.2 Why does self-feedback work?
Feedback capability refers to the capacity to provide information or guidance to a given input based on specific standards. As we discussed above, such capability is extensively utilized in various paradigms of automated alignment. For example, in RLAIF, LLM itself substitutes for human assessment to construct preference data (Yuan et al., 2024b; etc.). Additionally, LLM can continuously optimize its output based on self-feedback, a process known as self-refine (Bai et al., 2022b; Madaan et al., 2023; Tan et al., 2023; etc.). Nonetheless, there is much debate about whether and why LLMs can provide effective feedback to their own responses. In the following, we systematically summarize each representative viewpoint.
@@
Li et al. (2024g) and Lin et al. (2024b) suggest that models possess certain knowledge that cannot be directly utilized for generation but can be employed for providing feedback. Li et al. (2024g) also discover significant Generator-Validator inconsistency in LLMs when generating and validating answers. Similarly, Lin et al. (2024b) find that LLMs possess substantial knowledge that cannot be expressed through generation and correction but can be utilized for critique. Furthermore, other studies (Yuan et al., 2024b; Li et al., 2024f) argue that feedback capability is a byproduct of the model’s ability to follow instructions. Therefore, during the alignment process, the model’s feedback capacity improves alongside its ability to follow instructions, a phenomenon supported by experimental evidence (Yuan et al., 2024b; Li et al., 2024f). However, some other studies posit that the feedback capability of LLM is illusory, suggesting that it may rely on specific data and exhibit biases. West et al. (2023) note that while current LLMs gain generative capabilities to replicate expert-level outputs through training, they lack understanding capacities relevant to critique. Moreover, Huang et al. (2023b) observe that for reasoning tasks, LLMs even experience performance degradation after self-correction at times because that they cannot properly judge the correctness of their reasoning. Additionally, Zheng et al. (2024b) find that although powerful LLMs like GPT-4 achieve high agreement with human evaluators, similar to human-human agreement levels, LLM-based evaluation face various challenges such as position bias (Zheng et al., 2024b; Wang et al., 2023c), verbosity bias (Zheng et al., 2024b; Wu and Aji, 2023), self-enhancement bias (Zheng et al., 2024b; Liu et al., 2023b), as well as limited capability in certain scenarios such as grading math, reasoning questions (Zheng et al., 2024b), high-quality summaries (Shen et al., 2023a) and so on (Lan et al., 2024). Researchers propose many methods to further improve the capability of self-feedback, such as meta critic (Sun et al., 2024a), autocalibrate (Liu et al., 2023d), split and merge (Li et al., 2023e) and using external tools (such as search engines, code interpreters, etc.) to cross-check (Gou et al., 2024), thus refining their initial response, etc.

### v3

@@ 3.2.3 Discussion
Task decomposition and self-play both necessitate LLMs functioning as agents. Indeed, although research on agents is already very prosperous, the current capabilities of LLMs as agents are still very limited. They still struggle to complete tasks that would take humans several hours to finish (OpenAI, 2024). Therefore, focusing on improving the capabilities of LLMs-as-agents remains an important future research direction. This includes understanding and modeling the human objectives in the most economically valuable tasks, and developing effective reasoning abilities for ultra-long thought process.
Meanwhile, the challenge of aligning LLMs as agents is more complex compared to aligning them as chatbots, as it requires considerations of behavior-level alignment (Pan et al., 2023; Yuan et al., 2024b), the dynamics of environment and self-constraints (Garrabrant and Demski, 2018; Shavit et al., 2023; Yang et al., 2024f). More importantly, while adversarial self-play may enhance agent capabilities in long-horizon tasks, it can also give rise to the emergence of more deceptive (Hubinger et al., 2024), persuasive and autonomous agents (Tao et al., 2024). Such developments could have significant social impacts and ethical risks, such as the potential for models to generate more persuasive articles than humans, which could be exploited for political manipulation. Encouragingly, several prominent model engine providers have taken steps to monitor and mitigate these potential side effects. For example, OpenAI’s Preparedness team has established benchmarks for assessing persuasion and autonomy (OpenAI, 2023a), categorizing model risks into four levels and imposing development and deployment restrictions based on risk thresholds. Additionally, third-party organizations are contributing to the development of robust safety frameworks for highly capable agents (METR, 2024).
Furthermore, an even more intricate problem lies in proving the theoretical safety and trustworthiness of multi-agent systems. Although research in this area is nascent (Yang and Wang, 2020; DiGiovanni and Zell, 2021), advancements in game theory (Hazra and Anjaria, 2022), automated theorem proving techniques (Polu and Sutskever, 2020) and real-world simulation technology (Brooks et al., 2024) may offer insights into addressing this challenge.
@@ 5.1.3 Reinforcement Learning from AI Feedback
With the improved capabilities of LLMs, directly using them to rank multiple responses can provide preference data (Tunstall et al., 2023; Hong et al., 2023; Guo et al., 2024a; Pace et al., 2024; Yuan et al., 2024c). This ranked preference data can also be produced with minimal human supervision, such as through human-defined principles (Bai et al., 2022b; Sun et al., 2023c) or rules (Kim et al., 2023). To improve the quality of generated preference data, Shi et al. (2024) propose a carefully designed pipeline including reversed instruction tuning, instruction induction, and expert model evaluation.  Liu et al. (2024a) propose to score the response using contrastive prompt pairs as input, which can achieve better performance compared to directly generated feedback with a single prompt.
@@ 5.2 Binary Verifier
Even if the final answer is correct, there may still be errors in the reasoning process, limiting the effectiveness of the outcome verifier. To address this problem, the process verifier is employed to assess the correctness of reasoning step for more detailed verification (Lightman et al., 2023; Uesato et al., 2022). The process verifier can be used to train a more effective reasoner (Ying et al., 2024; Shao et al., 2024). Inspired by the human reasoning mechanism, Zhu et al. (2023b) propose Cooperative Reasoning (CoRe) to produce synthesized training data for reasoning where process verifier is used to generate the feedback of the generation of generator. Many studies are dedicated to training the verifier using automatically generated data due to the difficulty of collecting step-wise supervised signals.  Wang et al. (2024e) and Wang et al. (2024i) train the process verifier with automatically constructed data that is collected by Monte Carlo Sampling. Moreover, the process verifier can be applied in decoding to select the correct reasoning path (Khalifa et al., 2023). Some studies focus on how to complete the final reasoning path efficiently. Ma et al. (2023b) propose a heuristic greedy search algorithm based on the verifier’s feedback. Li et al. (2023c) propose using the verifier to filter the reasoning steps generated using diverse prompts.
@@ 6.2 Human Shared-Values
In order to achieve consensus on what basic principles AI systems should follow among groups of people and reach human shared-values by synthesizing various human ideas into a set of target for model alignment, follow-up efforts are dedicated to enable broader public to collectively shape the behavior of AI systems. For example, in a program titled Democratic inputs to AI333openai.com/blog/democratic-inputs-to-ai, an open call for the design of prototype systems to promote the democratic process of AI was issued. Specifically, it addresses enabling a representative cross-section of people to exchange views and ultimately reach a consensus on the rules AI systems should follow. Subsequently, in the Collective Constitutional AI (CCAI) project (Anthropic, 2023; Huang et al., 2024), researchers establish a multi-stage process for sourcing and integrating public input into LMs based on questionnaire. Through the way of collective constitution drafting, participants can assess the existing constitutional principles, rate their acceptance scores, and propose new constitutions they believe AI should be better to adhere to, thus reduce the bias of aligned LLMs. Beyond these pioneering research, large language models themselves are also involved to help conclude shared-values from human crowds more efficiently. Klingefjord et al. (2024) propose Moral Graph Elicitation (MGE) to elicit and reconcile values from human participants, which utilizes an LLM to interview participants about their values in particular contexts.
@@ 7 Underlying Mechanism of Automated Alignment
  • RQ2: Why does Self-feedback work? Self-feedback is widely applied across various paradigms of automated alignment, such as serving as an inductive bias in section 3.1.2, constructing data for preference learning in section 5 and acting as a crucial technique for automated iterative alignment (Yuan et al., 2024c; etc.). Understanding why does it work is vital for comprehending and enhancing all paradigms involving it.
@@ 7.2 Why does self-feedback work?
Feedback capability refers to the capacity to provide information or guidance to a given input based on specific standards. As we discussed above, such capability is extensively utilized in various paradigms of automated alignment. For example, in RLAIF, LLM itself substitutes for human assessment to construct preference data (Yuan et al., 2024c; etc.). Additionally, LLM can continuously optimize its output based on self-feedback, a process known as self-refine (Bai et al., 2022b; Madaan et al., 2023; Tan et al., 2023; etc.). Nonetheless, there is much debate about whether and why LLMs can provide effective feedback to their own responses. In the following, we systematically summarize each representative viewpoint.
@@
Li et al. (2024g) and Lin et al. (2024b) suggest that models possess certain knowledge that cannot be directly utilized for generation but can be employed for providing feedback. Li et al. (2024g) also discover significant Generator-Validator inconsistency in LLMs when generating and validating answers. Similarly, Lin et al. (2024b) find that LLMs possess substantial knowledge that cannot be expressed through generation and correction but can be utilized for critique. Furthermore, other studies (Yuan et al., 2024c; Li et al., 2024f) argue that feedback capability is a byproduct of the model’s ability to follow instructions. Therefore, during the alignment process, the model’s feedback capacity improves alongside its ability to follow instructions, a phenomenon supported by experimental evidence (Yuan et al., 2024c; Li et al., 2024f). However, some other studies posit that the feedback capability of LLM is illusory, suggesting that it may rely on specific data and exhibit biases. West et al. (2023) note that while current LLMs gain generative capabilities to replicate expert-level outputs through training, they lack understanding capacities relevant to critique. Moreover, Huang et al. (2023b) observe that for reasoning tasks, LLMs even experience performance degradation after self-correction at times because that they cannot properly judge the correctness of their reasoning. Additionally, Zheng et al. (2024b) find that although powerful LLMs like GPT-4 achieve high agreement with human evaluators, similar to human-human agreement levels, LLM-based evaluation face various challenges such as position bias (Zheng et al., 2024b; Wang et al., 2023c), verbosity bias (Zheng et al., 2024b; Wu and Aji, 2023), self-enhancement bias (Zheng et al., 2024b; Liu et al., 2023b), as well as limited capability in certain scenarios such as grading math, reasoning questions (Zheng et al., 2024b), high-quality summaries (Shen et al., 2023a) and so on (Lan et al., 2024). Researchers propose many methods to further improve the capability of self-feedback, such as meta critic (Sun et al., 2024a), autocalibrate (Liu et al., 2023d), split and merge (Li et al., 2023e) and using external tools (such as search engines, code interpreters, etc.) to cross-check (Gou et al., 2024), thus refining their initial response, etc.